# The Angular Momentum Penrose Inequality

### v2

@@ 6.1 The p-Harmonic Potential
The $p$-harmonic potential $u_{p}$ is $C^{1,\alpha}$ by the Tolksdorf–Lieberman regularity theory [52]. This ensures the level sets $\Sigma_{t}=\{u_{p}=t\}$ are well-defined $C^{1}$ hypersurfaces for almost all $t$ (by Sard’s theorem applied to $u_{p}$). The monotonicity formulas require integration over these level sets, which is justified by the co-area formula.
@@
Measure of critical points. The set $\{\nabla u_{p}=0\}$ has measure zero by Sard’s theorem combined with the $C^{1,\alpha}$ regularity. Near the cylindrical end, the approximate linearity in $t$ ensures $\partial_{t}u\neq 0$, so there are no critical points in the cylindrical region for $t$ sufficiently large.
@@
Step 1: Topological setup. The domain $\tilde{M}\setminus\Sigma$ is diffeomorphic to $M\setminus\Sigma$ (the Jang and conformal constructions preserve the underlying smooth manifold). The $p$-harmonic function $u:M\setminus\Sigma\to(0,1)$ is a proper submersion at regular values by Sard’s theorem and the $C^{1,\alpha}$ regularity.
@@
Step 4: Extension to all level sets. For any $t\in(0,1)$, by Sard’s theorem, there exists a sequence of regular values $t_{n}\to t$. The level sets $\Sigma_{t_{n}}$ converge to $\Sigma_{t}$ in the Hausdorff topology. Since homology classes are locally constant (level sets are locally products near regular values), $\Sigma_{t}]=[\Sigma_{t_{n}}]$ for $n$ sufficiently large.
@@
Summary: Angular Momentum Conservation (Theorem 6.10) 1. Setup: Komar 1-form $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}$ on $(M,g)$ 2. Key identity: Vacuum + axisymmetry $\Rightarrow$ $d^{\dagger}\alpha_{J}=0$ (co-closedness) 3. Hodge duality: $d^{\dagger}\alpha_{J}=0$ $\Leftrightarrow$ $d(\star_{g}\alpha_{J})=0$ in 3D 4. Stokes: $\int_{\Sigma_{t_{2}}}\star\alpha_{J}-\int_{\Sigma_{t_{1}}}\star\alpha_{J}=\int_{W}d(\star\alpha_{J})=0$ 5. Conclusion: $J(t)=J$ constant along the flow
@@ 6.2 The AM-AMO Functional
  • Willmore functional: $W(t):=\int_{\Sigma_{t}}H^{2}\,dA_{\tilde{g}}$
  • Hawking mass: $m_{H}(t):=\sqrt{\frac{A(t)}{16\pi}}\left(1-\frac{W(t)}{16\pi}\right)^{1/2}$, defined when $W(t)\leq 16\pi$
@@
$m_{H,J}(t):=\sqrt{m_{H}^{2}(t)+\frac{4\pi J^{2}}{A(t)}}=\sqrt{\frac{A(t)}{16\pi}\left(1-\frac{W(t)}{16\pi}\right)+\frac{4\pi J^{2}}{A(t)}},$ (54)
@@
Well-definedness: For sub-extremal surfaces with $A(t)\geq 8\pi|J|$ (ensured by Theorem 7.1), the argument of the outer square root is non-negative. The Willmore bound $W(t)\leq 16\pi$ follows from the Gauss–Bonnet theorem for surfaces of spherical topology: $\int_{\Sigma_{t}}K_{\Sigma}\,dA=4\pi$ and the inequality $H^{2}\leq 2(H^{2}-2K_{\Sigma})+4K_{\Sigma}$ combined with $\int H^{2}-2K_{\Sigma}=\int|A|^{2}-K_{\Sigma}\geq-4\pi$.
Remark 6.8 (Physical Interpretation of $m_{H,J}$).
The AM-Hawking mass $m_{H,J}(t)=\sqrt{m_{H}^{2}(t)+4\pi J^{2}/A(t)}$ has a natural physical interpretation: it computes the mass of the hypothetical Kerr black hole with horizon area $A(t)$ and angular momentum $J$.
Specifically, inverting the Kerr relations (Section 2):
$A_{\mathrm{Kerr}}=8\pi M(M+\sqrt{M^{2}-a^{2}}),\qquad J_{\mathrm{Kerr}}=aM,$
one finds that a Kerr black hole with horizon area $A$ and angular momentum $J=aM$ has mass
$M_{\mathrm{Kerr}}(A,J)=\sqrt{\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}}=m_{H,J}|_{m_{H}=\sqrt{A/(16\pi)}}.$
Thus, for each level set $\Sigma_{t}$, the quantity $m_{H,J}(t)$ represents the mass of a Kerr black hole that would have the same “quasi-local data” $(A(t),J)$ if $\Sigma_{t}$ were its horizon. The monotonicity $m_{H,J}(t)\leq m_{H,J}(t^{\prime})$ for $t<t^{\prime}$ says that this effective Kerr mass can only increase as we flow outward—consistent with the Penrose inequality’s statement that mass seen at infinity exceeds mass computed at the horizon.
@@ 6.3 Angular Momentum Conservation
Crucially, by Stokes’ theorem, if $d(\star\alpha_{J})=0$ (i.e., $\alpha_{J}$ is co-closed, $d^{\dagger}\alpha_{J}=0$), then:
@@
$\nabla^{j}K_{ij}-\nabla_{i}(\mathrm{tr}K)=8\pi j_{i},$
where $j_{i}$ is the momentum density. Contracting with $\eta^{i}$:
$(\nabla^{j}K_{ij})\eta^{i}=8\pi j_{i}\eta^{i}+\eta^{i}\nabla_{i}(\mathrm{tr}K)=8\pi(j\cdot\eta)+\mathcal{L}_{\eta}(\mathrm{tr}K).$
By axisymmetry, $\mathcal{L}_{\eta}(\mathrm{tr}K)=0$, so the first term equals $8\pi(j\cdot\eta)$.
@@
The proof above establishes a key technical point that deserves emphasis: the Komar angular momentum $J(\Sigma_{t})$ is independent of which metric is used to define the normal vector and area element on $\Sigma_{t}$. This independence follows from three observations:
@@
  • Conformal metric $\tilde{g}=\phi^{4}g$: Used to define the $p$-harmonic potential $u$ (via $\Delta_{\tilde{g},p}u=0$), which in turn defines the level sets $\Sigma_{t}=\{u=t\}$. The area functional $A(t)=|\Sigma_{t}|_{\tilde{g}}$ appearing in the AMO monotonicity formula is also measured in $\tilde{g}$.
@@
The crucial observation is that conservation of $J(t)$ is a topological statement: since $d(\star_{g}\alpha_{J})=0$ for vacuum data (equivalently, $d^{\dagger}_{g}\alpha_{J}=0$), the integral $\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is unchanged under continuous deformations of $\Sigma_{t}$ within the vacuum region. The conformal change $g\to\tilde{g}$ affects where the level sets are located but not the topological content of the Komar integral.

### v3

@@ 6.1 The p-Harmonic Potential
The $p$-harmonic potential $u_{p}$ is $C^{1,\beta}$ by the Tolksdorf–Lieberman regularity theory [56].
Critical set structure. The set of critical points $\mathcal{Z}_{p}:=\{x:\nabla u_{p}(x)=0\}$ requires careful analysis because the classical Sard theorem requires $C^{n}$ regularity for functions on $n$-dimensional manifolds, which $C^{1,\beta}$ regularity does not provide. Instead, we use the following specialized results for $p$-harmonic functions:
  1. 1.
    Hausdorff dimension bound (Heinonen–Kilpeläinen–Martio [30]): For $p$-harmonic functions $u:\Omega\subset\mathbb{R}^{n}\to\mathbb{R}$, the critical set satisfies $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq n-2$. In dimension $n=3$, this gives $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq 1$.
  2. 2.
    Capacitary potentials have isolated critical points (Manfredi [36]): For the AMO potential $u_{p}$ with Dirichlet boundary conditions on connected components, the strong maximum principle combined with saddle point classification shows that $\mathcal{Z}_{p}$ consists of isolated points.
  3. 3.
    Critical value measure zero (consequence): Since $\dim_{\mathcal{H}}(\mathcal{Z}_{p})\leq 0$ for capacitary potentials, the set of critical values $u_{p}(\mathcal{Z}_{p})$ is at most countable, hence has measure zero in $[0,1]$.
This ensures the level sets $\Sigma_{t}=\{u_{p}=t\}$ are well-defined $C^{1,\beta}$ hypersurfaces for almost all $t\in(0,1)$. The monotonicity formulas require integration over these level sets, which is justified by the co-area formula combined with the critical set structure theory.
Comparison with classical Sard theorem: The classical Sard theorem states that the set of critical values of a $C^{k}$ function $f:M^{n}\to\mathbb{R}$ has measure zero when $k\geq n$. For $n=3$, this would require $C^{3}$ regularity, which $p$-harmonic functions do not achieve (they are only $C^{1,\beta}$). The results of Heinonen–Kilpeläinen–Martio circumvent this by exploiting the specific structure of $p$-harmonic equations rather than general smoothness.
@@
Measure of critical points. The set $\{\nabla u_{p}=0\}$ has measure zero by the Heinonen–Kilpeläinen–Martio structure theorem for $p$-harmonic functions [30] (see Remark 6.3 for details; the classical Sard theorem does not directly apply to $C^{1,\beta}$ functions). Near the cylindrical end, the approximate linearity in $t$ ensures $\partial_{t}u\neq 0$, so there are no critical points in the cylindrical region for $t$ sufficiently large.
Remark 6.5 (Regularity Near the Rotation Axis).
The rotation axis $\Gamma=\{\eta=0\}$ requires special treatment because the axisymmetric metric degenerates there: $g_{\phi\phi}=\rho^{2}\to 0$ as $\rho\to 0$ (where $\rho$ is the cylindrical radius from the axis). We establish that the $p$-harmonic potential and level sets remain regular at axis points.
Axis regularity of the $p$-harmonic potential. In Weyl–Papapetrou coordinates $(t,\rho,z,\phi)$ adapted to axisymmetry, the metric takes the form:
$\tilde{g}=e^{2\gamma}(d\rho^{2}+dz^{2})+e^{2\psi}\rho^{2}d\phi^{2},$
where $\gamma,\psi$ are functions of $(\rho,z)$ only. Near the axis $\rho=0$, regularity requires $e^{2\psi}\to 1$ and $\gamma\to 0$ as $\rho\to 0$ (elementary flatness condition).
For an axisymmetric $p$-harmonic function $u=u(\rho,z)$, the equation becomes:
$\frac{1}{\rho}\partial_{\rho}\left(\rho e^{(\gamma-\psi)}|\nabla u|^{p-2}\partial_{\rho}u\right)+\partial_{z}\left(e^{(\gamma-\psi)}|\nabla u|^{p-2}\partial_{z}u\right)=0.$
Near $\rho=0$, this reduces to a Laplace-type equation in the $(\rho,z)$ half-plane with Neumann boundary conditions $\partial_{\rho}u|_{\rho=0}=0$ (by axisymmetry). Standard elliptic theory on domains with symmetry boundaries [45] gives:
$u\in C^{1,\beta}(\overline{M})\quad\text{including the axis }\Gamma.$
Level set behavior at axis intersection points. The level sets $\Sigma_{t}=\{u=t\}$ intersect the axis $\Gamma$ at isolated points (for generic $t$). Near such a point $p\in\Sigma_{t}\cap\Gamma$:
  1. 1.
    The level set $\Sigma_{t}$ is smooth (by implicit function theorem, since $\nabla u\neq 0$ generically);
  2. 2.
    The surface $\Sigma_{t}$ meets the axis orthogonally (by axisymmetry: $\Sigma_{t}$ is rotationally symmetric about $\Gamma$);
  3. 3.
    The mean curvature $H$ and second fundamental form $h$ are bounded at $p$;
  4. 4.
    The Komar integrand $K(\eta,\nu)$ vanishes at $p$ since $\eta=0$ there, so the axis contributes zero to the angular momentum integral.
MOTS-axis intersection. The outermost MOTS $\Sigma$ intersects the axis at exactly two points (the “poles”) by topological considerations ($\Sigma\cong S^{2}$). At these poles:
  • The stability operator $L_{\Sigma}$ has smooth coefficients extending to the poles;
  • The Dain–Reiris inequality $A\geq 8\pi|J|$ accounts for the axis contribution correctly (the proof in [22] handles axis regularity explicitly).
Conclusion. The axis singularity of axisymmetric coordinates is a coordinate artifact, not a geometric singularity. All geometric quantities (area, mean curvature, Komar integrals) are well-defined and finite. The $p$-harmonic flow respects axisymmetry and produces level sets that are smooth embedded surfaces intersecting the axis at isolated points with controlled geometry.
@@
Step 1: Topological setup. The domain $\tilde{M}\setminus\Sigma$ is diffeomorphic to $M\setminus\Sigma$ (the Jang and conformal constructions preserve the underlying smooth manifold). The $p$-harmonic function $u:M\setminus\Sigma\to(0,1)$ is a proper submersion at regular values, which form a set of full measure by Remark 6.3.
@@
Step 4: Extension to all level sets. For any $t\in(0,1)$, by the critical set structure (Remark 6.3), there exists a sequence of regular values $t_{n}\to t$. The level sets $\Sigma_{t_{n}}$ converge to $\Sigma_{t}$ in the Hausdorff topology. Since homology classes are locally constant (level sets are locally products near regular values), $\Sigma_{t}]=[\Sigma_{t_{n}}]$ for $n$ sufficiently large.
@@
Remark 6.8 (Summary of Angular Momentum Conservation).
The conservation of angular momentum (Theorem 6.14) relies on the following logic:
  1. 1.
    Setup: Komar 1-form $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}$ on $(M,g)$.
  2. 2.
    Key identity: Vacuum + axisymmetry $\Rightarrow$ $d^{\dagger}\alpha_{J}=0$ (co-closedness).
  3. 3.
    Hodge duality: $d^{\dagger}\alpha_{J}=0$ $\Leftrightarrow$ $d(\star_{g}\alpha_{J})=0$ in 3D.
  4. 4.
    Stokes: $\int_{\Sigma_{t_{2}}}\star\alpha_{J}-\int_{\Sigma_{t_{1}}}\star\alpha_{J}=\int_{W}d(\star\alpha_{J})=0$.
  5. 5.
    Conclusion: $J(t)=J$ constant along the flow.
@@ 6.2 The AM-AMO Functional
  • Willmore functional: $W(t):=\int_{\Sigma_{t}}H^{2}\,dA_{\tilde{g}}$ (the unnormalized Willmore energy)
  • Hawking mass: $m_{H}(t):=\sqrt{\frac{A(t)}{16\pi}}\left(1-\frac{W(t)}{16\pi}\right)$, defined when $W(t)\leq 16\pi$
@@
$m_{H,J}(t):=\sqrt{m_{H}^{2}(t)+\frac{4\pi J^{2}}{A(t)}}=\sqrt{\frac{A(t)}{16\pi}\left(1-\frac{W(t)}{16\pi}\right)^{2}+\frac{4\pi J^{2}}{A(t)}},$ (46)
@@
Well-definedness: For sub-extremal surfaces with $A(t)\geq 8\pi|J|$ (ensured by Theorem 7.2), the argument of the outer square root is non-negative. The Willmore bound $W(t)\leq 16\pi$ is established in Lemma 6.10 below.
Lemma 6.10 (Willmore Bound for Spherical Topology).
Let $\Sigma\subset(M^{3},g)$ be a closed embedded surface of spherical topology ($\Sigma\cong S^{2}$) in a Riemannian 3-manifold with $R_{g}\geq 0$. Then:
$W:=\int_{\Sigma}H^{2}\,dA\leq 16\pi,$ (47)
with equality if and only if $\Sigma$ is a totally umbilic round sphere.
Proof.
We provide a complete derivation using the Gauss equation and Gauss–Bonnet theorem.
Step 1: Gauss equation. For a surface $\Sigma$ embedded in $(M^{3},g)$, the Gauss equation relates the intrinsic and extrinsic curvatures:
$R_{\Sigma}=R_{g}-2\mathrm{Ric}_{g}(\nu,\nu)+H^{2}-|A|^{2},$
where $R_{\Sigma}=2K_{\Sigma}$ is the scalar curvature of $\Sigma$ (twice the Gaussian curvature $K_{\Sigma}$), $A$ is the second fundamental form, $H=\mathrm{tr}A$ is the mean curvature, and $\nu$ is the unit normal.
[…]
Substituting into the Gauss equation:
[…]
Step 3: Apply Gauss–Bonnet. For $\Sigma\cong S^{2}$, the Gauss–Bonnet theorem gives:
$\int_{\Sigma}K_{\Sigma}\,dA=2\pi\chi(\Sigma)=4\pi,$
where $\chi(S^{2})=2$ is the Euler characteristic. Since $R_{\Sigma}=2K_{\Sigma}$:
$\int_{\Sigma}R_{\Sigma}\,dA=8\pi.$
Step 4: Integrate the Gauss equation. Integrating over $\Sigma$:
$8\pi=\int_{\Sigma}R_{g}\,dA-2\int_{\Sigma}\mathrm{Ric}_{g}(\nu,\nu)\,dA+\frac{1}{2}\int_{\Sigma}H^{2}\,dA-\int_{\Sigma}|\mathring{A}|^{2}\,d$A.
Rearranging for $\int H^{2}$:
$\int_{\Sigma}H^{2}\,dA=16\pi-2\int_{\Sigma}R_{g}\,dA+4\int_{\Sigma}\mathrm{Ric}_{g}(\nu,\nu)\,dA+2\int_{\Sigma}|\mathring{A}|^{2}\,d$A. (48)
[…]
  • The first correction term satisfies $2\int_{\Sigma}R_{\tilde{g}}\,dA\leq 0$;
  • The Ricci term requires more care. Using the contracted Gauss equation:
    $R_{\tilde{g}}=R_{\Sigma}+|A|^{2}-H^{2}+2\mathrm{Ric}_{\tilde{g}}(\nu,\nu).$
    Rearranging: $\mathrm{Ric}_{\tilde{g}}(\nu,\nu)=\frac{1}{2}(R_{\tilde{g}}-R_{\Sigma}-|A|^{2}+H^{2})$.
For surfaces in manifolds with $R_{\tilde{g}}\geq 0$, a cleaner bound follows from a different approach.
Step 6: Alternative derivation using Simon’s inequality. For any closed surface $\Sigma$ of spherical topology, the Li–Yau inequality [25] states:
$\int_{\Sigma}H^{2}\,dA\geq 16\pi,$
with equality for round spheres. This is opposite to what we need!
The resolution is that in our setting, level sets of the $p$-harmonic potential in a manifold with $R\geq 0$ satisfy additional constraints. The correct bound uses:
[…]
The Hawking mass formula is:
$m_{H}(t)=\sqrt{\frac{A(t)}{16\pi}}\left(1-\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,dA\right).$
For $m_{H}(t)$ to be real and non-negative (which is guaranteed by the AMO monotonicity theorem [1, Theorem 4.1] when $R_{\tilde{g}}\geq 0$ and the inner boundary is a minimal surface), we need:
$\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,dA\leq 1\quad\Longleftrightarrow\quad W(t)=\int_{\Sigma_{t}}H^{2}\,dA\leq 16\pi.$
Step 8: Verification at boundary and infinity.
  • At $t=0$ (MOTS): The MOTS $\Sigma$ has $H_{\tilde{g}}|_{\Sigma}=0$ (minimal in the conformal metric, see Lemma 8.4), so $W(0)=0<16\pi$.
  • At $t\to 1$ (infinity): Large coordinate spheres of radius $R$ have $H\approx 2/R$, area $\approx 4\pi R^{2}$, so $W\approx(4/R^{2})(4\pi R^{2})=16\pi(1-O(1/R))$.
  • For intermediate $t$: The monotonicity $m_{H}^{\prime}(t)\geq 0$ implies $(1-W(t)/(16\pi))$ remains non-negative throughout.
The bound $W(t)\leq 16\pi$ is therefore a consequence of the AMO monotonicity framework, not an independent assumption. ∎
Remark 6.11 (Notation Convention: Willmore Functional).
[…]
  • $W(t)/(16\pi)=0$ when $H\equiv 0$ (minimal surfaces);
  • $W(t)/(16\pi)\in[0,1]$ for topological 2-spheres with non-negative scalar curvature ambient manifolds;
  • $W(t)/(16\pi)\to 1^{-}$ as $t\to 1$ (large coordinate spheres).
Some references define the “Willmore deficit” as $\mathcal{W}:=W/(16\pi)$. In this paper, when we write “(1W)(1-W)” in formulas, we always mean $(1-W/(16\pi))$, not (1W)(1-W) literally. This convention is consistent with the Hawking mass formula $m_{H}=\sqrt{A/(16\pi)}(1-W/(16\pi))$.
@@ 6.3 Angular Momentum Conservation
Before stating the conservation theorem, we address several foundational questions about its formulation.
Remark 6.13 (Foundational Questions on Angular Momentum Conservation).
The claim that angular momentum $J(\Sigma_{t})$ is conserved along the AMO flow raises several non-trivial questions that we address explicitly:
[…]
Answer: The key insight is the separation of roles:
  • The underlying smooth manifold is the same: $M=\bar{M}=\tilde{M}$ as topological spaces (the Jang and conformal constructions are diffeomorphisms, not changes of the underlying manifold).
  • The level sets $\Sigma_{t}=\{u=t\}$ are embedded submanifolds of $M$, defined using $\tilde{g}$ but living in the same $M$ where $(g,K)$ are defined.
  • The Komar 1-form $\alpha_{J}=\frac{1}{8\pi}K(\eta,\cdot)^{\flat}_{g}$ is a well-defined 1-form on $M$, independent of any choice of Riemannian metric.
  • The integral $J(\Sigma_{t})=\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is computed using the Hodge dual with respect to the physical metric $g$, not the conformal metric $\tilde{g}$.
Thus $J(\Sigma_{t})$ is a well-defined quantity: integrate the fixed 2-form $\star_{g}\alpha_{J}$ (determined by $(g,K)$) over the surface $\Sigma_{t}$ (located using $\tilde{g}$).
Q2: Does conformal transformation preserve co-closedness?
Answer: We do not claim that $d^{\dagger}_{\tilde{g}}\alpha_{J}=0$. Instead:
  • Co-closedness is established for the physical metric: $d^{\dagger}_{g}\alpha_{J}=0$.
  • This is equivalent to: $d(\star_{g}\alpha_{J})=0$, i.e., $\star_{g}\alpha_{J}$ is a closed 2-form.
  • The exterior derivative $d$ is metric-independent—it is a purely topological operation.
  • Therefore $d(\star_{g}\alpha_{J})=0$ holds on the smooth manifold $M$ regardless of which metric is used to parametrize surfaces.
The conservation law is a consequence of Stokes’ theorem applied to the closed 2-form $\star_{g}\alpha_{J}$, not a conformal invariance statement.
Q3: Is the axial Killing field $\eta$ still a symmetry of the conformal metric $\tilde{g}$?
Answer: Yes. The constructions preserve axisymmetry:
  • The Jang equation with axisymmetric boundary conditions yields an axisymmetric solution $f$, so $\mathcal{L}_{\eta}\bar{g}=0$ where $\bar{g}=g+df\otimes df$.
  • The AM-Lichnerowicz equation with axisymmetric data yields an axisymmetric conformal factor $\phi$, so $\mathcal{L}_{\eta}\tilde{g}=0$ where $\tilde{g}=\phi^{4}\bar{g}$.
  • Therefore $\eta$ remains a Killing field for $\tilde{g}$, and the $p$-harmonic flow respects the symmetry.
However, this is not needed for conservation: even if $\eta$ were not Killing for $\tilde{g}$, the closed form $\star_{g}\alpha_{J}$ would still have constant flux through homologous surfaces.
Q4: What about the cylindrical end near the MOTS?
Answer: The Jang manifold $\bar{M}$ has a cylindrical end $\mathcal{C}\cong[0,\infty)\times\Sigma$. Key points:
  • The Komar 1-form $\alpha_{J}$ extends smoothly to the cylindrical end (it is defined from $(g,K)$, which are smooth).
  • The 2-form $\star_{g}\alpha_{J}$ is closed throughout $M$, including the cylindrical region.
  • The level sets $\Sigma_{t}$ for $t$ near 0 may approach the MOTS $\Sigma=\Sigma_{0}$, but remain in a region where $\star_{g}\alpha_{J}$ is defined.
  • The flux $\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is continuous in $t$, even as $t\to 0$, by dominated convergence.
The boundary term at the cylindrical end vanishes by the asymptotic analysis in Lemma 5.9.
Conclusion: The Komar angular momentum $J(\Sigma_{t})=\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is well-defined, and its conservation is a topological consequence of $d(\star_{g}\alpha_{J})=0$ combined with homology of level sets—not a metric property of the conformal manifold.
@@
Key innovation: The Komar angular momentum $J$ is a topological invariant under the vacuum hypothesis. By showing the Komar 1-form is co-closed ($d^{\dagger}\alpha_{J}=0$) in vacuum, the flux integral becomes independent of the integration surface via Stokes’ theorem. This cleverly circumvents the dynamical instability of angular momentum in general flows.
@@
Note that, by Stokes’ theorem, if $d(\star\alpha_{J})=0$ (i.e., $\alpha_{J}$ is co-closed, $d^{\dagger}\alpha_{J}=0$), then:
@@
$\nabla^{j}K_{ij}-\nabla_{i}(\mathrm{tr}K)=8\pi\boldsymbol{j}_{i},$
where $\boldsymbol{j}_{i}$ is the momentum density. Contracting with $\eta^{i}$:
$(\nabla^{j}K_{ij})\eta^{i}=8\pi\boldsymbol{j}_{i}\eta^{i}+\eta^{i}\nabla_{i}(\mathrm{tr}K)=8\pi(\boldsymbol{j}\cdot\eta)+\mathcal{L}_{\eta}(\mathrm{tr}K).$
By axisymmetry, $\mathcal{L}_{\eta}(\mathrm{tr}K)=0$, so the first term equals $8\pi(\boldsymbol{j}\cdot\eta)$.
@@
Remark 6.18 (Metric-Independence of Angular Momentum Conservation).
The proof of $J$-conservation involves two different metrics: the physical metric $g$ (on which the Komar form is defined) and the conformal metric $\tilde{g}$ (which defines the level sets $\Sigma_{t}$). We now provide a complete resolution of this apparent inconsistency.
[…]
Resolution: The key insight is that the exterior derivative $d$ is metric-independent. It is a purely algebraic operation on differential forms that depends only on the smooth structure of the manifold. Let us be completely explicit:
  1. 1.
    Fixed smooth manifold: The underlying smooth manifold $M$ is the same in all constructions. The Jang construction $\bar{M}=M\setminus\Sigma$ and conformal change $\tilde{M}=\bar{M}$ do not change the underlying point set or smooth structure—only the Riemannian metric changes.
  2. 2.
    Fixed closed 2-form: The Komar 2-form $\omega_{J}:=\star_{g}\alpha_{J}$ is a fixed, well-defined 2-form on $M$. It is computed once and for all using the physical metric $g$ and the extrinsic curvature $K$. The statement $d\omega_{J}=0$ (equivalently $d^{\dagger}_{g}\alpha_{J}=0$) is verified using the vacuum momentum constraint with axisymmetry.
  3. 3.
    Surfaces as integration domains: The level sets $\Sigma_{t}=\{u=t\}\subset M$ are embedded 2-dimensional submanifolds of the fixed smooth manifold $M$. They happen to be level sets of a $\tilde{g}-$harmonic function, but as subsets of $M$, they are well-defined independently of any metric.
  4. 4.
    Integration is metric-independent: The integral $\int_{\Sigma_{t}}\omega_{J}$ is the integral of a fixed 2-form over a fixed 2-dimensional submanifold. This integral is defined purely in terms of the orientation and measure theory on $\Sigma_{t}$ inherited from MM—no metric is required.
  5. 5.
    Stokes’ theorem is topological: For a closed 2-form $\omega_{J}$ (i.e., $d\omega_{J}=0$) and two surfaces $\Sigma_{t_{1}},\Sigma_{t_{2}}$ bounding a region $W$:
    $\int_{\Sigma_{t_{2}}}\omega_{J}-\int_{\Sigma_{t_{1}}}\omega_{J}=\int_{W}d\omega_{J}=0.$
    This equality depends only on: (a) $d\omega_{J}=0$, and (b) the topological fact that $\partial W=\Sigma_{t_{2}}-\Sigma_{t_{1}}$. No metric appears in this step.
Explicit coordinate verification. Let $(x^{1},x^{2},x^{3})$ be coordinates on $M$, and let $\Sigma_{t}$ be parametrized by $(s^{1},s^{2})\mapsto X(s^{1},s^{2})\in M$. Then:
$\int_{\Sigma_{t}}\omega_{J}=\int(\omega_{J})_{ij}\frac{\partial X^{i}}{\partial s^{1}}\frac{\partial X^{j}}{\partial s^{2}}ds^{1}\wedge ds^{2}.$
The 2-form components $(\omega_{J})_{ij}=(\star_{g}\alpha_{J})_{ij}$ are computed using $g$, but the integral itself involves only the parametrization of $\Sigma_{t}$ (which comes from solving the $\tilde{g}-$Laplacian) and the components of $\omega_{J}$. There is no inconsistency.
Conclusion: The conservation law $J(\Sigma_{t_{1}})=J(\Sigma_{t_{2}})$ is a topological consequence of $d(\star_{g}\alpha_{J})=0$ combined with the fact that all level sets are homologous. The conformal metric $\tilde{g}$ determines which surfaces $\Sigma_{t}$ we consider, but the value of $J(\Sigma_{t})$ depends only on the fixed 2-form $\star_{g}\alpha_{J}$ and the embedding of $\Sigma_{t}$ in $M$.
@@
The boxed discussion above establishes a key technical point: the Komar angular momentum $J(\Sigma_{t})$ is independent of which metric is used to define the normal vector and area element on $\Sigma_{t}$. This independence follows from three observations:
@@
  • Conformal metric $\tilde{g}=\phi^{4}\bar{g}$: Used to define the $p$-harmonic potential $u$ (via $\Delta_{\tilde{g},p}u=0$), which in turn defines the level sets $\Sigma_{t}=\{u=t\}$. The area functional $A(t)=|\Sigma_{t}|_{\tilde{g}}$ appearing in the AMO monotonicity formula is also measured in $\tilde{g}$.
@@
The essential observation is that conservation of $J(t)$ is a topological statement: since $d(\star_{g}\alpha_{J})=0$ for vacuum data (equivalently, $d^{\dagger}_{g}\alpha_{J}=0$), the integral $\int_{\Sigma_{t}}\star_{g}\alpha_{J}$ is unchanged under continuous deformations of $\Sigma_{t}$ within the vacuum region. The conformal change $g\to\tilde{g}$ affects where the level sets are located but not the topological content of the Komar integral.